OpenStudy (anonymous):

solution=-1 Apply the quadratic formula.

OpenStudy (anonymous):

the quadratic formula can only be applied to quadratic functions written in the form: ax^2 + bx + c = o where a, b, c are all real numbers.

OpenStudy (radar):

The quadratic equation whose solution is -1, if factored would be (x+1)(x+1)=0 Working that out would be: x^2+2x+1=0